Find the width (in meters) of a rectangle with perimeter 15.4 meters and length 3.4 meters.
Answer #1

Solutions length ( l ) = Boum width (w ]=w Perimeter = 2(lt) = 15.4 => 2 (3.47 w)=15.4 → 3.4+ W = 7.7 - W= 7.7-3.4 = 4.3 the